OpenStudy (anonymous):

please how do you solve this equation sin3x+2cos3x=o

OpenStudy (anonymous):

what are you looking for?

OpenStudy (anonymous):

for values of x

OpenStudy (anonymous):

idk if this is what your looking for but i got x= (tan^-1(-2))/3

OpenStudy (aravindg):

u r from which country folk?

OpenStudy (anonymous):

Mauritius Island

OpenStudy (anonymous):

ok fernandoruiz thx u

OpenStudy (anonymous):

There are two real solutions from the Mathematica Reduce function: \[x=\frac{1}{3} \left(2 \text{ArcTan}\left[\frac{1}{2} \left(1-\sqrt{5}\right)\right]+2 \pi C[1]\right) \]and\[x=\frac{1}{3} \left(2 \text{ArcTan}\left[\frac{1}{2} \left(1+\sqrt{5}\right)\right]+2 \pi C[1]\right) \]where C[1] is an integer. The solutions were verified using C[1] = 0.

OpenStudy (vishweshshrimali5):

the solutions are : \[x = \frac{2}{3}(\pi . n + tan^{-1}(\frac{1}{2}(1-\sqrt {5})))\] and \[x = \frac{2}{3}(\pi . n + tan^{-1}(\frac{1}{2}(1+\sqrt {5})))\]

OpenStudy (anonymous):

great work vishweshshrimali5 bhaiya